# Deconstructing Obfuscation: A four-dimensional framework for evaluating Large Language Models assembly code deobfuscation capabilities

### v1

@@ 4.3 Limitations
Our study focused on seven commercial LLMs representing current state-of-the-art capabilities but cannot capture the full landscape. Notably absent are domain-specific code models and open-source alternatives that might exhibit different capabilities. This limitation reflects the practical constraints of comprehensive evaluation rather than intentional exclusion. Future research should evaluate specialized code-oriented models that may demonstrate different dimensional strengths.
@@ 5 Conclusion
Our comprehensive evaluation of state-of-the-art LLMs on assembly-level deobfuscation establishes a new paradigm for understanding AI capabilities in cybersecurity contexts. This research delivers a robust theoretical framework, decisive evidence, and far-reaching implications that transform our understanding of automated code analysis.
The results demonstrate clear performance patterns across the obfuscation techniques. Several models deobfuscated bogus control flow with minimal intervention, while all models - regardless of architecture or size - failed against combined obfuscation techniques. Control flow flattening and instruction substitution presented intermediate challenges, with performance varying significantly across models. This performance gradient definitively maps the boundaries of current LLM capabilities for advanced binary analysis tasks.
The four-dimensional framework we introduce: Reasoning Depth, Pattern Recognition, Noise Filtering, and Context Integration provide essential explanatory mechanisms for these variations. Each obfuscation technique targets specific dimensional capabilities: bogus control flow tests reasoning depth through invariant conditions; instruction substitution challenges pattern recognition through equivalent operations; control flow flattening disrupts context integration by fragmenting related code; and combined techniques simultaneously undermine all dimensions. This mapping conclusively explains the inconsistent performance across techniques and establishes a theoretical foundation that advances our understanding of LLM deobfuscation capabilities.

### v2

@@ 4.3 Limitations
Architectural Scope
Our evaluation focused exclusively on x86_64 assembly, which may limit the generalizability of our dimensional framework to other instruction set architectures. While x86_64 represents a dominant platform for desktop and server environments, embedded systems and mobile devices increasingly employ ARM, RISC-V, and other architectures that may present different deobfuscation challenges.
@@
Our study focused on eight commercial LLMs representing current state-of-the-art capabilities but cannot capture the full landscape. Notably absent are domain-specific code models and open-source alternatives that might exhibit different capabilities. This limitation reflects the practical constraints of comprehensive evaluation rather than intentional exclusion. Future research should evaluate specialized code-oriented models that may demonstrate different dimensional strengths.
@@ 5 Conclusion
Our comprehensive evaluation of state-of-the-art LLMs on assembly-level deobfuscation establishes a new paradigm for understanding AI capabilities in cybersecurity contexts. This research delivers a robust theoretical framework, compelling evidence, and far-reaching implications that transform our understanding of automated code analysis.
The results demonstrate clear performance patterns across the obfuscation techniques. Several models deobfuscated bogus control flow with minimal intervention, while all models - regardless of architecture or size - failed against combined obfuscation techniques. Control flow flattening presented minimal challenge while instruction substitution proved highly resistant, with performance varying significantly across models. This performance gradient definitively maps the boundaries of current LLM capabilities for advanced binary analysis tasks.
The four-dimensional framework we introduce: Reasoning Depth, Pattern Recognition, Noise Filtering, and Context Integration provide essential explanatory mechanisms for these variations. Each obfuscation technique targets specific dimensional capabilities: bogus control flow tests reasoning depth through invariant conditions; instruction substitution challenges pattern recognition through equivalent operations; control flow flattening disrupts context integration by fragmenting related code; and combined techniques simultaneously undermine all dimensions. This mapping helps explain the inconsistent performance across techniques and establishes a theoretical foundation that advances our understanding of LLM deobfuscation capabilities.